OpenStudy (anonymous):

How many pounds of cashews selling for $7.40 per pound must be mixed with 6 pound of peanut selling for $2.60 to create a mixture that sells for $5.80 per pound?

OpenStudy (anonymous):

7.40(x) + 2.60(6) = 5.80(x + 6)

OpenStudy (anonymous):

7.40x + 15.60 = 5.80x + 34.80

OpenStudy (anonymous):

1.60x = 19.20

OpenStudy (anonymous):

x = 12 12 pounds of cashews

OpenStudy (anonymous):

thanks blexting